OpenStudy (anonymous):

HELPPP!! Solve this equation algebraically - 2^x+2^-x=5

OpenStudy (anonymous):

2^x + 1/2^x = 5 (2^x)^2 + 1 = 5 (2^x) ??

OpenStudy (anonymous):

(2^x)= ???

OpenStudy (anonymous):

\[2^x + 2^{-x} = 5 = 2^x + \frac{1}{2^x}\] \[(5) 2^x = 2^x 2^x + 1\]

OpenStudy (anonymous):

\[\log 5 + x(\log 2) = 2x(\log 2) + \log(1) = 2x(\log 2)\]

OpenStudy (anonymous):

thank youuu